OpenStudy (anonymous):

Find the graphing form, coordinates of the center & radius, then graph the points and equation. X2 + 2x + y2 – 6y – 71 = 0

OpenStudy (anonymous):

@sourav_aich

OpenStudy (anonymous):

the graphing form is a circle x^2 + 2x +1 -1 y^2 + 2.3.y + 9 -9 -71 = 0 (x+1)^2 + (y+3)^2 - 81 = 0 (x+1)^2 + (y+3)^2 = 9^2 so centre is (-1,-3) and radius is 9 units